the segment joining A(-3, "-2)" and B(-1/2 , 3) is extended to a distance equal to 1/5 of its own length. find the new end point closest to point A
Answer by greenestamps(13327)



It is implied in the wording of the problem that one of the points remains stationary. So if we want to find the point closest to A after increasing the length of AB by 1/5 of its length, we need to move from point A AWAY FROM point B.

The difference between A and B in the x direction is 2.5 units (the difference between -3 and -1/2); the difference in the y direction is 5 (the difference between -2 and 3).

Extending AB by 1/5 of its length means adding 1/5 of 2.5 = 0.5 units to the difference in the x direction and 1/5 of 5 = 1 unit to the difference in the y direction.

B is to the right and above A; to keep B stationary and increase the length of the segment by 1/5 of its length, we need to start at A and move 0.5 units LEFT and 1 unit DOWN -- putting us at (-3-0.5,-2-1) = (-3.5,-3).

ANSWER: (-3.5,-3)